The following table gives the joint probability distribution between employment status & college graduation in the...

  1. The following table gives the joint probability distribution between employment status & college graduation in the US working age population for 2012.
    Unemployed (Y=0) Employed (Y=1) Total
    Non-college graduates (X=0) 0.053 0.586 0.639
    College graduates (X=1) 0.015 0.346 0.361
    Total 0.068 0.932 1
    • Calculate the unemployment rate for college graduates, & for non-college graduates.
    • A randomly selected member of the population reports being unemployed. What is the probability that they are a college graduate? A non-college graduate?
    • Are educational achievement & employment status independent? Explain.
    • Does this relationship prove that college education has a causal effect on employment status? Use the potential outcomes framework to explain why or why not.
Answer #1

Sokution un employed Employed! Toł al ty-1) (Y=0) Non-Colleoge gadustes Cx =0) 0.6 39 O.053 O. 586 0. 346 College Jrakuotes (in dependany d). then. P(M,y) = Let Hen P (0,0) = 0.053 P(x-0) = O-639 P Cy=0) = 0.068. Na P(X=0) . pcy=0) - o.639x0.068 = 0.